OpenStudy (anonymous):

Given the function g(x) = −2x + 4, find g(−5).

OpenStudy (jim766):

substitute the -5 where ever you see an x

OpenStudy (jim766):

so it would be g(-5) = -2(-5) + 4 then simplify

OpenStudy (jim766):

can you do it from here?

OpenStudy (anonymous):

-5g=-10+4 divide all by 5 right?

OpenStudy (anonymous):

g(x) is a function.. You cannot write it as : xg You have to write in g(x) form.. So when you plug x = -5 here: It will remain like this: \(g(-5)\)..

OpenStudy (anonymous):

Can you tell me the value for this: \(-2(-5) = ??\)

OpenStudy (anonymous):

my bad... 10

OpenStudy (anonymous):

So what will be the value now can you tell ??

OpenStudy (anonymous):

\[g(-5) = -2(-5) + 4 = ??\]

OpenStudy (anonymous):

-5g=10+4

OpenStudy (anonymous):

I said to leave that in g(-5) form.. You will not do anything with left hand side.. g(-5) will remain as it is.. Why are you not adding 10 and 4 ???

OpenStudy (anonymous):

ohh ok and idk i added them together in my head.

OpenStudy (anonymous):

So what is the value for \(g(-5)\) ??

OpenStudy (anonymous):

wouldn't it be g(-5)?

OpenStudy (anonymous):

Yes.. What I have written ??

OpenStudy (anonymous):

g(-5)=14 so how would d i solve this?

OpenStudy (anonymous):

You have solved this.. This is what question wants from you.. g(-5) = 14 is the answer dear..

OpenStudy (anonymous):

well now it makes sense... i didn't understand the question.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Welcome..